Prepare the variance analysis for the following and show formulas in words and work

Prepare One Month Answer Significance (e.g., "Favorable" or "Unfavorable")

Groomer Direct Labor Time Variance

DL Time Variance = (AH - SH) x SR _______________ _________________

Groomer Directv Labor Rate Variance

DL Rate Variance = (AR - SR) x AH ________________ __________________

Direct Materials Efficiency

DM Quantity Variance = (AQ - SQ) xSP ________________ __________________

Direct Materials Price Variance

DM Price Variance = (AP - SP) x AQ ________________ __________________

For your variance analysis, use the following financial data:

Grooming labor Standard Hours Standard Rate Actual Hours Actual Rate

150 $12 180 $11.50

Grooming Materials Standard Quantity Standard Price Actual Quantity Actual Price

1000 $2 1200 $3
Answer #1

Groomer Direct labour Time variance = (AH-SH)*SR

= (180-150)*12

Direct labour time variance = 360 Unfavorable

Groomer Direct labour rate variance = (AR-SR)*AH

= (11.50-12)*180

Groomer direct labour rate variance = 90 Favorable

Direct material quantity variance = (AQ-SQ)*SP

= (1200-1000)*2

Direct material quantity variance = 400 Unfavorable

Direct material price variance = (AP-SP)*AQ

= (3-2)*1200

Direct material price variance = 1200 Unfavorable